OpenStudy (anonymous):

determine the domain of the function: fx=sqrt 3-x

OpenStudy (anonymous):

(-infinity, 3)

OpenStudy (zarkon):

\[(-\infty,3]\]

OpenStudy (anonymous):

\[3<x\]

OpenStudy (zarkon):

no...\[x\leq 3\]

OpenStudy (zarkon):

\[\sqrt{3-x}\geq 0\] \[3-x\geq 0\] \[3\geq x\]